OpenStudy (anonymous):

HELP I'LL GIVE A METAL!! Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. cot x sec4x = cot x + 2 tan x + tan3x

OpenStudy (anonymous):

\cot x \sec^4 x = \cot x + 2 \tan x + \tan^3 x

OpenStudy (anonymous):

cotx+2tanx+tan3x 1tanx+2tanx+tan3x 1+2tan2x+tan4xtanx 1+tan2x+tan2x+tan4xtanx sec2x+tan2x+tan4xtanx sec2x+tan2x(1+tan2x)tanx sec2x+tan2x(sec2x)tanx sec2x(1+tan2x)tanx sec2x(sec2x)tanx sec4xtanx cotxsec4x = LEFT HAND SIDE

OpenStudy (anonymous):

That's it?

OpenStudy (welshfella):

i dont get some of that

OpenStudy (anonymous):

sec^2=tna^1+1

OpenStudy (welshfella):

cot x sec4x = cos x 1 1 ---- * --- = ------- sin x cos^4 x sinx cos^3 x

OpenStudy (welshfella):

now i've managed to prove that RHS = same thing:-

OpenStudy (welshfella):

changing RHS to sin and cos = cos x + 2 sin x + sin^3 x ----- ----- ------- sin x cos x cos^3 x = cos^4 x + 2 sin^2 x cos ^2 x + sin^4x --------------------------------- sin x cos^3 x = (cos^2 x + sin^2 x)(cos^2 x + sin^2 x) -------------------------------- sin x cos^3 x = 1 * 1 ----- = LHS sin x cos^3 x

OpenStudy (anonymous):

Ok... Your legit.

OpenStudy (anonymous):

Is that it?

OpenStudy (welshfella):

yes both LHS and RHS = 1 /sin x cos^3 x so they must be equal

OpenStudy (welshfella):

its a bit long winded and there might be an easier bit it is legit.

OpenStudy (anonymous):

Same question, different problem. (sin x)(tan x cos x - cot x cos x) = 1 - 2 cos2x

OpenStudy (welshfella):

you know the identity cos^2 x + sin^2 x = 1 right?

OpenStudy (welshfella):

I dont get GoArmy's solution . It might be right but i'm not sure...

OpenStudy (welshfella):

similar way convert LHS to sin and cos = sin x( sin x cos x - cos x cos x) ---- --- cos x sin x = sin x( sin x - cos^2 x) ------- sin x = sin x( sin^2 x - cos^2 x) ---------------- sin x = sin^2 x - cos^2 x

OpenStudy (welshfella):

now can you finish this off?

OpenStudy (welshfella):

use the identity sin^2x + cos^2 x = 1

OpenStudy (welshfella):

substitute for sin^2 x